07 February 2008

RHIC pp Set-Up:  

Spin pattern checked at Fill and it is OK

Measured large emittances (50pi) due to quad polarity reversal in AtR.  Fixed.

Feedback ramps got to 99% efficiency.

Nonlinear chromaticity correction

Attempted feedback ramps using store tunes.  Failed twice and abandoned.  Back to tune swing ramps.

Measured polarization at store.  36% in Blue,10% in Yellow.

Rotator ramp set up during evening shift.  Second rotator ramp gave 97% blue and 89% yellow.

Blue horizontal store emittance at 20pi.  Wrong Blue spin pattern.

Provided collisions over night with 37x37.  7.5kHz at PHENIX, 120kHz BBC at both experiments.

JET set-up during store.

Tested 10Hz IR feedbacks.  BPM signals noisy; setpoints saturate.

Yellow horizontal store emittance 32pi, depolarized.  Next store gave 18pi with 15% polarization.

Blue decay still shows one minute periodicity.

Continue ramp tuning including polarization..  Do feedback ramps, nonlinear chromaticity correction.  IR nonlinear correction.  Check 10Hz feedback system..  Increase intensity and number of bunches.

There is an impending LINAC failure that needs a 1-2 hour access behind a store. 

STAR requests an access for ~30 minutes to fix a stuck trigger bit.  Standing request for access by PP2PP.  PHENIX has requested a one hour access some time this week in order to fix a TDC.

PHENIX needs to make measurements with vertical polarization to calibrate their detectors.  Then another ramp with radial polarization for the remainder of the run.

Rob Michnoff reports that he observed the same kind of timing shifts between d-Au and pp (ie 1ns at PHENIX and 17ns at STAR) and he suspects that there was a phase shift in the V124 module providing the timing pulse.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020